OpenStudy (anonymous):

Help? D: How do you determine the exact trigonometric ratio of tan 5pi/12?

OpenStudy (eyust707):

well 5/12 = 3/12 + 2/12 right?

OpenStudy (anonymous):

..Yeah

OpenStudy (eyust707):

and 3/12 is 1/4 2/12 is 1/6

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

I understand that far.. Kind of

OpenStudy (eyust707):

sweet so that means that: \[\tan ({5\pi \over 12} )= \tan ({3\pi \over 12} + {2\pi \over 12}) = \tan ({\pi \over 4} + {\pi \over 6})\]

OpenStudy (anonymous):

That makes sense so far, it's usually what follows that ends up looking like Chinese. >.<

OpenStudy (eyust707):

Well then you can use an identity...

OpenStudy (eyust707):

the bottom one should work

OpenStudy (anonymous):

Do you mind showing me how it would work out?.. >.< Sorry

OpenStudy (eyust707):

No its fine =) as long as you help me

OpenStudy (eyust707):

what would you use for u?

OpenStudy (eyust707):

what about v?

OpenStudy (anonymous):

so u can be 1/4 and v can be 1/6

OpenStudy (eyust707):

Yep! make sure you use (1/4)* pi\[\tan ({\pi \over 4} + {\pi \over 6})\]

OpenStudy (eyust707):

u = \(\pi \over 4\) v = \(\pi \over 6\)

OpenStudy (eyust707):

we end up with